# PAID: A Framework of Product-Centric Advertising Image Design

### v2

@@ 4.5. Graphics Rendering
Since there is no effective method to predict the visual attributes of graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering. We first sort the tagline elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024b). The selection is based on the similarity to the foreground color and contrast with the background area. Taglines grouped by size and position use the same font and color. For the underlay, the color is chosen based on its contrast with the tagline color and similarity to the background color. The shape is selected from a predefined Scalable Vector Graphics (SVG) library based on its size and proportions, with minor adjustments for fitting.
@@ 5. Experiments
In this section, we compare our PAID framework and its modules with previous ones. We conducted experiments on the PITA dataset, and the results confirm the effectiveness of our method.
@@ 5.1. Implementation Details
For prompt and layout generation, XCP2 (Dong et al., 2024) is used as our backbone. Its visual encoder is built upon CLIP ViT-Large (Radford et al., 2021), processing 490x490 resolution images. The LLM is based on InternLM2 (Cai et al., 2024). We fine-tune the LLM branch on the PITA dataset for one epoch. We train the model using AdamW optimizer with a learning rate of 1e-5. Batch size on each GPU is set to 1, and gradient accumulation of 8 is adopted. The training costs about 14 hours with 8 NVIDIA A100 GPUs. During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
For background image generation, we use the SDXL base model as the backbone. During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used. We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution. Then in step 2, we introduce an inpainting ControlNet, which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
@@ 5.2. Evaluation Metrics
To evaluate the overall poster generation and each module, we use the following metrics.
Overall pipeline. We evaluate the overall pipeline from three perspectives: overall visual quality, and layout. For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score. The aesthetic score comes from advertising experts ranking the outputs of different methods. The average rank represents the mean aesthetic rating, with a lower rank indicating better creative results. For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
Prompt generation. We evaluate prompt generation quality based on fore-background matching rate (FBM rate) and e-commerce domain score (ED score). FBM Rate refers to how well a product foreground fits with a given prompt background. We classify this matching into “reasonable” or “unreasonable” through human annotation, and calculate the proportion of reasonable cases. ED Score assesses how well the model-generated prompts reflect e-commerce characteristics. We calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
@@
Layout generation. Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic metrics and content metrics to evaluate layouts. Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$. Val means the ratio of elements larger than 0.1% of the canvas. Ali assesses how well elements are spatially aligned. Ove is the average intersection over union (IoU) of all element pairs except for underlays. $Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements. Content metrics assess the harmony of layout with the image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$. Uti evaluates space usage for graphic elements. Occ is the average overlapping area between the graphic elements and products. Rea represents the non-flatness of regions that tagline elements without underlays are put on. For methods that take taglines as input, we report the tagline match rate (TMR), to see if the number of generated tagline bboxes is consistent with input taglines.
Background image generation. We evaluate the quality of generated images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2022). To verify layout control, we use Grounding-DINO (Liu et al., 2024c) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth bboxes. The average of these maximum IoUs is called mIOU. In addition, Grounding-DINO AP scores are also reported.
@@ 5.3. Comparison
Compare with automatic pipelines for advertising image design. Since there are no public methods for advertising image design conditioned on the product foreground, taglines, and target size, we design two pipelines for comparison. Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation. We also finetune the SDXL-based inpainting model on our PIL dataset. The only difference between this method and ours is the absence of layout control and the corresponding training stage. During inference, only the product foreground is controlled by the layout output of GPT-4o. Pipeline 2 first generate a background prompt using GPT-4o. The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general aesthetic rules. Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout. Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image. Table 1 shows the quantitative comparison. PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box. However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality. PAID uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 5). Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals. Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded tagline boxes, making them hard to read, while our model provides more flexible and readable layouts.
@@
PAID 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
Compare with prompt generation methods. For prompt comparison, we use GPT-4o and an untrained version of our model (XCP2) as baselines. Both utilize in-context learning (ICL) for optimal performance. The FBM Rate is determined by votes from three annotators for each case. As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score. This means that our model effectively aligns with the e-commerce prompt dataset distribution after training. We also visualize the feature distribution of prompts generated in the appendix.
@@
Compare with layout generation methods. To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a). For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased/generated set. The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased areas. The generated set contains images re-generated using their ground-truth prompts (image captions). The generated set has no above hints and conforms to real applications. As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and foreground. Our method uses the ground-truth prompt, product foreground, and taglines for prediction. The results are shown in Table 3 and Figure 6. Background-based methods perform relatively poorly in terms of product occlusion and overlap and alignment of graphic elements. This may be due to the fixed background image content, leaving less room for the model flexibility. RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating tagline boxes that cannot adequately fill the tagline content. The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images. Our method, PAID, takes into account the shape of the product foreground, background prompt, and tagline content to decide the overall layout, including graphic and nongraphic elements, resulting in a better overall effect.
@@
Compare with layout-controlled inpainting models. We compare our layout-controlled background image generation model with others, including training-based method GLIGEN (Li et al., 2023c) and training-free methods BoxDiff (Xie et al., 2023), TFCLG (Chen et al., 2024), and MultiDiffusion (Bar-Tal et al., 2023). Since these methods lack inpainting ability, we re-implement them with the SDXL-based inpainting ControlNet (Creative, 2024). Quantitative results are reported in Table 4. Training-free methods have weaker spatial control compared to training-based methods. Our trained model outperforms others in FID, mIoU, and AP, indicating better image quality and spatial control. As shown in Figure 7, other models tend to miss some instances when more objects are required, while our model generates them more accurately.
Figure 7. Visualization of images generated by different layout-controlled inpainting models.
Table 4. Quantitative performance comparison of layout-controlled inpainting models on PIL test set.
@@ 5.4. Online Result
To assess the online performance of PAID, we conduct A/B tests in two advertising recommendation scenarios on Taobao. We randomly select 5,000 products from various categories (including clothing, cosmetics, food, and electronics) and collect their foreground images and taglines from the advertisers. We then generate advertising images for each product using both PAID and Pipeline 2 (the top-performing baseline as shown in Table 1). The A/B tests are conducted using 5% of the main traffic, affecting only the experimental products when this traffic is directed to them. After gathering data over one month, we found that PAID achieved a 3.02% and 3.03% increase in click-through rate (CTR) for the two recommendation scenarios, respectively. This indicates that PAID is capable of producing more visually appealing advertising images, resulting in improved recommendation outcomes.
@@ 6. Conclusion
In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create advertising images only based on product foreground images, marketing taglines, and target sizes. PAID consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering. First, it generates adaptive prompts according to the product foregrounds. Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout. Next, a layout-controlled inpainting model is utilized for background image generation. A graphics rendering module is then applied to get the final images. We train separate experts to conduct these sub-tasks. Two datasets are created for the convenience of training and testing. The results of the test set and the online A/B tests have proven that the proposed PAID framework generates more visually pleasing and attractive advertising images.
@@ Acknowledgements
This work is supported by National Natural Science Foundation of China (NSFC No.62076031 and No.62076036) and Alibaba Research Intern Program.
@@ Appendix A Dataset Examples
In this section, we provide examples from the PITA and PIL datasets, illustrated in Figures 8 and 9, respectively. The PITA dataset includes advertising images with marketing taglines. In addition to depicting the layout of both graphic and nongraphic elements as shown in the figures, we also label the image caption (prompt), tagline content, and product foreground masks to train the prompt and layout generation models. Conversely, the PIL dataset contains product images without taglines. We annotate the layout of nongraphic elements and product foreground masks for training the image generation model.

### v3

@@ 4.5. Graphics Rendering
Since there is no effective method to predict the visual attributes of graphic elements, and the graphics rendering module is not the focus of this paper, we design a rendering strategy. We sort taglines by area to select the font and color from a limited set (Liu et al., 2024b). The choice depends on the similarity to the foreground color and contrast with the background. Taglines grouped by size and position use the same font and color. The underlay color is picked for contrast with the tagline and similarity to the background, while the shape is selected from a predefined SVG library based on size and proportions, with slight modifications to fit.
@@ 5. Experiments
In this section, we compare our framework with previous models.
@@ 5.1. Implementation Details
We use XCP2 (Dong et al., 2024) as the backbone for prompt and layout generation, fine-tuning the LLM branch on the PITA dataset for one epoch respectively. The model is trained with the AdamW optimizer at a 1e-5 learning rate, using a total batch size of 64. Training takes 14 hours on 16 NVIDIA H20 GPUs. For inference, we apply top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
For background image generation, we use the SDXL model as the backbone. In step 1, we train new layers using 6.8 million aesthetic images (from the LAION (Schuhmann et al., 2022) and PIL dataset) with layout annotations at a learning rate of 5e-5 for 20k steps on 512x512 resolution, taking 65 hours on 16 NVIDIA H20 GPUs (batch size 64). We then fine-tune them with a learning rate of 2e-5 for 20k steps at 1024x1024 resolution over 86 hours (batch size 32). In step 2, we introduce an inpainting ControlNet pre-trained on advertising images (Creative, 2024). We fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps using 12k training data from the PIL dataset, which takes 5 hours with 16 NVIDIA H20 GPUs (batch size 16).
@@ 5.2. Evaluation Metrics
We use the following metrics to assess the pipeline and each module.
Overall pipeline. We assess the pipeline from two perspectives: visual quality and layout. For visual quality, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) and an aesthetic score, with the latter based on rankings by advertising experts. A lower average rank indicates better results. For layout evaluation, we use metrics from previous work (Zhou et al., 2022; Seol et al., 2024), as detailed in the layout generation module.
Prompt generation. We evaluate prompt generation quality based on fore-background matching rate (FBM rate) and e-commerce domain score (ED score). FBM Rate measures how well a product foreground fits with the background prompt. We classify this matching into “reasonable” or “unreasonable” through human annotation, and calculate the proportion of reasonable cases. ED Score assesses how well the model-generated prompts reflect e-commerce characteristics. We calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
@@
Layout generation. Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic/content metrics to evaluate layouts. Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$. Content metrics assess the harmony of layout with the image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$. For methods that take taglines as input, we report the tagline match rate (TMR), to see if the number of generated tagline bboxes matches the input taglines.
Background image generation. We evaluate the quality of generated images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2022). To verify layout control, we use Grounding-DINO (Liu et al., 2024c) to detect instances and calculate the mIOU between detected bboxes and ground-truth bboxes. In addition, Grounding-DINO AP scores are also reported.
@@ 5.3. Comparison
Compare with automatic pipelines for advertising image design. Since there are no public methods for advertising image design conditioned on the product foregrounds, taglines, and target sizes, we design two pipelines for comparison. Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation. It fine-tunes an SDXL-based inpainting model on our PIL dataset. The main difference from our approach is the lack of layout control. During inference, only the product foreground is controlled by the layout output of GPT-4o. Pipeline 2 first generate a background prompt using GPT-4o. The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general aesthetic rules. Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout. Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image. Table 1 shows the quantitative comparison. T-Stars-Poster outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box. However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality. T-Stars-Poster uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 5). Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas T-Stars-Poster adjusts layouts based on product shape for better visuals. Additionally, the fixed background of Pipeline 2 can crowd tagline boxes, reducing readability, while our model generates more flexible and readable layouts.
@@
T-Stars-Poster 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
Compare with prompt generation methods. For prompt comparison, we use GPT-4o and an untrained version of our model (XCP2) as baselines. Both utilize in-context learning (ICL) for optimal performance. The FBM Rate is determined by votes from three annotators for each case. As shown in Table 2, our method outperforms the baselines, which means that our model effectively aligns with the e-commerce prompt dataset distribution after training.
@@
Compare with layout generation methods. We compare our methods with background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method, P&R (Li et al., 2023a). For background-based methods, we test on two sets for fair comparison: the erased set, with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on erased areas; and the generated set, which contains re-generated images with their ground-truth prompts (image captions), aligning with real applications. Since P&R has not released the codes, we re-implemented its idea. Ground-truth prompts are also given to P&R and our method. The results are shown in Table 3 and Figure 6. Background-based methods struggle with product occlusion, overlap, and alignment of graphic elements, likely due to fixed background content. RADM and DS-GAN also have issues with tagline boxes that cannot properly fit the content. P&R improves in these areas but sometimes generates unreasonable images due to a lack of harmony between product, background prompt, and graphic layout. T-Stars-Poster considers product shape, background prompt, and tagline content, optimizing the entire layout for graphic and non-graphic elements, ultimately achieving better overall results.
@@
Compare with layout-controlled inpainting models. We compare our layout-controlled background image generation model with others: the training-based method GLIGEN (Li et al., 2023c) and training-free methods like BoxDiff (Xie et al., 2023), TFCLG (Chen et al., 2024), and MultiDiffusion (Bar-Tal et al., 2023). Since these methods lack inpainting ability, we re-implement them with the SDXL-based inpainting ControlNet (Creative, 2024). The results are reported in Table 4. Our trained model outperforms others in FID, mIoU, and AP, indicating better image quality and spatial control.
Table 4. Quantitative comparison of layout-controlled inpainting models on PIL test set.
@@
5.4. User-Centric Study
To thoroughly evaluate the visual effects and attractiveness of the generated results, we randomly select 200 test samples and ask 7 Taobao users to compare images in pairs, evaluating background aesthetics, layout harmony, text readability, and overall visual appeal. As shown in Table 5, our method significantly outperforms the comparison methods in all aspects.
Table 5. User-centric study. The values in the table represent the winning rate of our method compared to other methods.
Comparison Method Bg Aesthetic Layout Harmony Text Readability Visual Appeal
Pipeline 1 84% 93% 94% 91%
Pipeline 2 77% 85% 89% 74%
@@ 5.5. Online Result
To evaluate the online performance of T-Stars-Poster, we conduct A/B tests in two advertising recommendation scenarios on Taobao. We randomly select 5,000 products and collect their foreground images and taglines from advertisers. We generate images for each product using both T-Stars-Poster and Pipeline 2 (the top-performing baseline as shown in Table 1). The A/B tests use 5% of the main traffic, affecting only the experimental products when this traffic is directed to them. After gathering data over one month, results show that T-Stars-Poster achieved a 3.02% and 3.03% increase in click-through rate (CTR) in the two recommendation scenarios, respectively. This indicates that T-Stars-Poster produces more visually appealing advertising images, resulting in improved recommendation outcomes.
6. Limitations and Discussion
Through observation, we have found that the results may be unsatisfactory in some cases. 1) Difficulty in adjusting product angle and lighting. As shown in Figure 5, inpainting models cannot adjust the product angle and lighting due to the model structure and training data, possibly causing disharmony in lighting and composition. While subject-driven methods (Tan et al., 2024) may help, they currently cannot fully preserve product features. 2) Limited flexibility with multiple foregrounds. As Figure 5 shows, multiple foregrounds are treated as one entity. This limits adjustments when products are spaced apart or angled very differently. Future improvements could involve treating separate subjects individually and introducing position augmentations during training. 3) Poor layout with many taglines. Over five taglines increase overlap and affect readability due to limited training data (4.3%) and canvas space. More data with multiple taglines might help resolve this.
@@ 7. Conclusion
This paper introduces T-Stars-Poster, a product-centric framework for creating advertising images using product foregrounds, taglines, and target sizes. It comprises four stages: prompt generation, layout generation, background image generation, and graphics rendering. First, it generates prompts according to the product foregrounds. Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout. Next, a layout-controlled inpainting model is utilized for background image generation. A graphics rendering module is applied to get the final images. We train separate experts to conduct these sub-tasks. Two datasets are created for the convenience of training and testing. Test results and online A/B tests show that T-Stars-Poster produces more visually pleasing and attractive advertising images.
@@ Acknowledgements
This work is supported by Alibaba Research Intern Program.
@@ Appendix A Dataset Examples
In this section, we provide examples from the PITA and PIL datasets, illustrated in Figures 7 and 8, respectively. The PITA dataset includes advertising images with marketing taglines. In addition to depicting the layout of both graphic and nongraphic elements as shown in the figures, we also label the image caption (prompt), tagline content, and product foreground masks to train the prompt and layout generation models. Conversely, the PIL dataset contains product images without taglines. We annotate the layout of nongraphic elements and product foreground masks for training the image generation model.